# Representations of categories of finite relational structures and associated endomorphism monoids

### 18 Mar 2026 v1

@@ 7.5. Application II
Let $\mathscr{E}\leqslant\operatorname{End}(\Omega,\mathcal{R}_{G})$ and $\mathscr{F}$ a dense submonoid of $\mathscr{E}$. Then
@@
We show that $\mathscr{E}$ and $\mathscr{F}$ have the same associated category $\mathscr{C}$} of finite substructures. For any finite substructure $A\subseteq\Omega$, we must verify that
@@
By the proof, we see that a submonoid $\mathscr{F}\subseteq\mathscr{E}$ induces the same associated category $\mathscr{C}$} of finite substructures as $\mathscr{E}$ if and only if $\mathscr{F}$ is dense in $\mathscr{E}$ with respect to the pointwise convergence topology, equivalently, if every finite restriction of an element of $\mathscr{E}$ is realized by an element of $\mathscr{F}$.
@@
Let $\mathscr{E}\leqslant\operatorname{End}(\Omega,\mathcal{R}_{G})$ and $\mathscr{F}$ a dense submonoid of $\mathscr{E}$. Then a continuous representation $V$ of $\mathscr{E}$ is uniformly continuous if and only if its restriction to $\mathscr{F}$ is uniformly continuous.
@@
Let $\mathscr{E}=\operatorname{End}(\Omega,\mathcal{R}_{G})$ be the full endomorphism monoid, and $\mathscr{E}_{i}$ be the submonoid of injective endomorphisms. Note that finite restrictions of elements in $\mathscr{E}_{i}$ are injective, and every injective morphism between finite substructures can be extended to an automorphism in $G$ by the homogeneity. Therefore, $G$ is dense in $\mathscr{E}_{i}$, so they have the same uniformly continuous representation theory. Furthermore, since continuous representations of $G$ are uniformly continuous, by Lemma 7.14, every continuous representation of $\mathscr{E}_{i}$ is uniformly continuous.
For instance, the monoid of injective maps in $\Omega^{\Omega}$, or the monoid of injective order-preserving maps in $\mathbb{Q}^{\mathbb{Q}}$ has the same continuous representation theory as $G=\operatorname{Sym}(\Omega)$ or $\operatorname{Aut}(\mathbb{Q},\leqslant)$, which has been studied in [10].
In contrast, the submonoid $\mathscr{E}_{s}$ of surjective endomorphisms is in general not dense in $\mathscr{E}$, but it is true when the action of $G$ on $\Omega$ is highly homogeneous. For instance, if $G=\operatorname{Sym}(\Omega)$, then $\mathscr{E}_{s}$ is dense in $\mathscr{E}=\Omega^{\Omega}$: given any map $f:A\to B$ between finite subsets, there exists a surjection $\tilde{f}\in\mathscr{E}_{s}$ such that $\tilde{f}|_{A}=f$. The reader can verify similar conclusion for the other four cases. Consequently, for highly homogeneous relational structures, $\mathscr{E}$ and $\mathscr{E}_{s}$ have the same uniformly continuous representation theory.
To consider uniform continuity of continuous representations of $\mathscr{E}=\operatorname{End}(\Omega,\mathcal{R}_{G})$, we introduce the following concept.
@@
We say that $(\Omega,\mathcal{R})$ satisfies the finite retraction property if for every finite subset $A\subseteq\Omega$, there exists an endomorphism $h\in\mathscr{E}=\operatorname{End}(\Omega,\mathcal{R})$ such that
@@
For instance, the endomorphism monoid $\mathscr{E}=\Omega^{\Omega}$ clearly satisfies the finite retraction property.
@@
If $(\Omega,\mathcal{R})$ satisfies the finite retraction property, then every continuous representation of $\mathscr{E}$ is uniformly continuous.
@@
The dense linear order $(\mathbb{Q},\leqslant)$ does not satisfy the finite retraction property. Indeed, for any finite subset $A\subseteq\mathbb{Q}$ with at least two points, there is no endomorphism $h\in\mathscr{E}=\operatorname{End}(\mathbb{Q},\leqslant)$ such that $h|_{A}=\operatorname{id}_{A}$ and $\mathrm{im}(h)\subseteq A$, because any order-preserving map from a dense linear order into a finite set must be locally constant, and thus cannot fix a set with more than one point. Consequently, $\mathscr{E}$ admits continuous representations that are not uniformly continuous.
@@ 7.6. Application III
For the third application, we let $\mathscr{E}\leqslant\operatorname{End}(\Omega,\mathcal{R}_{G})$ and assume that the associated category $\mathscr{C}$} of finite substructures satisfies the following condition
@@
    every inclusion $\iota_{A,B}:A\hookrightarrow B$ admits a retraction given by a surjective morphism $p:B\twoheadrightarrow A$.
The five categories associated to the endomorphism monoids of highly homogeneous relational structures satisfy this condition. Besides, there are many other examples.
Example 7.20.
Let $G=\operatorname{Sym}(\Omega)$ act diagonally on the set $\Omega\times\Omega$. This action is not highly homogeneous. However, it is easy to verify that an injective morphism $\iota:A\hookrightarrow B$ in $\mathscr{C}$} admits a retraction $p:B\twoheadrightarrow A$ blockwise, so $\mathscr{C}$} satisfies $(\sharp)$ as well.
@@
Let $\mathscr{E}$ be a submonoid of $\operatorname{End}(\Omega,\mathcal{R}_{G})$, and suppose that the associated category $\mathscr{C}$} of finite substructures satisfies condition $(\sharp)$. Then the realization functor $\mathcal{R}$ gives an equivalence
@@
Let $A$ be a finite substructure and let $S\in J(A)$ be a covering sieve on ${\mathscr{C}}^{{\mathrm{op}}}$, which is the same as a left ideal on $\mathscr{C}$}. By the description of the Grothendieck topology $J$, the sieve $S$ contains an inclusion $\iota_{A,B}:A\hookrightarrow B$. By condition $(\sharp)$, there is a surjective morphism $p:B\twoheadrightarrow A$ such that
@@
Let $(\Omega,\mathcal{R}_{G})$ be a highly homogeneous relational structure. Without loss of generality, we assume that $\Omega=\mathbb{Q}$, on which the five highly homogeneous relational structures have been explicitly described in [2, Section 13]. The equivalence described in Corollary 7.21 provides us with a useful approach to study uniformly continuous representations of $\mathscr{E}=\operatorname{End}(\mathbb{Q},\mathcal{R}_{G})$.
@@
Let $k$ be a commutative ring and fix $n\in\mathbb{Z}_{+}$. Let $G_{n}$ be the automorphism group of the finite substructure $[n]$. Then under the equivalence in Corollary 7.21, we have the following correspondences:
@@
If furthermore $k$ is a field of characteristic 0 and $\lambda\in\operatorname{Irr}(G_{n})$, then
$(k\mathscr{M}_{n}/k\mathscr{N}_{n})\otimes_{kG_{n}}\lambda\,\longleftrightarrow\,\Delta_{n,\lambda}.$
@@
One can then use the above result to obtain an explicit construction of irreducible uniformly continuous $k\mathscr{E}$-modules when $k$ is a field of characteristic 0.
@@
Let $k$ be a field of characteristic 0. The isomorphism classes of irreducible uniformly continuous $k\mathscr{E}$-modules are parameterized by the set of pairs
@@
Specifically, let $V_{n,\lambda}=(k\mathscr{M}_{n}/k\mathscr{N}_{n})\otimes_{kG_{n}}\lambda$. Then the irreducible $k\mathscr{E}$-modules are described as follows:
@@
    If $\lambda$ is regular, the corresponding irreducible $k\mathscr{E}$-module is isomorphic to $V_{n,\lambda}$.
@@
    If $\lambda$ is singular, the corresponding irreducible $k\mathscr{E}$-module is isomorphic to the top of $V_{n,\lambda}$.
@@ 7.7. Application IV
The proofs of Proposition 7.6, Theorem 7.11, and Corollary 7.21 apply in this setting with suitable modifications. The orbit relations on $\Omega$ induced by the action of $G=\mathrm{GL}(\Omega)$ encode linear dependence of tuples. Consequently, the relevant finite substructures are precisely the finite dimensional $\mathbb{F}_{q}$-subspaces of $\Omega$. It follows that the category constructed from restrictions of elements of $\mathscr{E}$ is precisely ${\mathrm{VA}}_{q}$. Moreover, if $f,g\in\mathscr{E}$ coincide on a finite subset $A\subseteq\Omega$, then they coincide on the $\mathbb{F}_{q}$-linear span $\langle A\rangle$, since $f$ and $g$ are linear maps. Thus uniformly continuous $k\mathscr{E}$-modules can be defined via finite dimensional subspaces. Putting these observations together, we obtain the first equivalence in the first line.

### 25 Apr 2026 v3

@@ 7.5. Application II
Let $\mathscr{E}\leqslant\Omega^{\Omega}$ and $\mathscr{F}$ a dense submonoid of $\mathscr{E}$. Then
@@
We show that $\mathscr{E}$ and $\mathscr{F}$ have the same associated category $\mathscr{C}$} of finite subsets. For any finite subset $A\subseteq\Omega$, we must verify that
@@
By the proof, we see that a submonoid $\mathscr{F}\subseteq\mathscr{E}$ induces the same associated category $\mathscr{C}$} of finite subsets as $\mathscr{E}$ if and only if $\mathscr{F}$ is dense in $\mathscr{E}$ with respect to the pointwise convergence topology, equivalently, if every finite restriction of an element of $\mathscr{E}$ is realized by an element of $\mathscr{F}$.
@@
Let $\mathscr{E}\leqslant\Omega^{\Omega}$ and $\mathscr{F}$ a dense submonoid of $\mathscr{E}$. Then a continuous representation $V$ of $\mathscr{E}$ is uniformly continuous if and only if its restriction to $\mathscr{F}$ is uniformly continuous.
@@
Let $\mathscr{E}=\Omega^{\Omega}$ be the full transformation monoid, and let $\mathscr{E}_{i}$ be the submonoid of injective maps. Note that finite restrictions of elements of $\mathscr{E}_{i}$ are injective, and every injective map between finite subsets extends to a permutation in $G=\operatorname{Sym}(\Omega)$. It follows that $G$ is dense in $\mathscr{E}_{i}$ with respect to the permutation topology, so $\mathscr{E}_{i}$ and $G$ have the same uniformly continuous representation theory. Moreover, since continuous representations of $G$ are automatically uniformly continuous, Lemma 7.14 implies that every continuous representation of $\mathscr{E}_{i}$ is uniformly continuous.
Similarly, let $\mathscr{E}=\operatorname{End}(\mathbb{Q},\leqslant)$ be the monoid of order-preserving maps. Then the submonoid $\mathscr{E}_{i}$ of injective order-preserving maps has the same uniformly continuous representation theory as $\operatorname{Aut}(\mathbb{Q},\leqslant)$, which has been studied in [12]. The same conclusion holds for other highly homogeneous relations on $\mathbb{Q}$, namely the dense cyclic order, dense linear betweenness, and dense separation relations.
On the other hand, the submonoid $\mathscr{E}_{s}$ of surjective maps is also dense in $\mathscr{E}=\Omega^{\Omega}$. Indeed, given any map $f:A\to B$ between finite subsets, there exists a surjective map $\tilde{f}\in\mathscr{E}_{s}$ such that $\tilde{f}|_{A}=f$. Analogous density statements hold for the corresponding monoids of maps preserving a dense linear order, cyclic order, betweenness relation, or separation relation on $\Omega=\mathbb{Q}$.
To consider uniform continuity of continuous representations of $\mathscr{E}$, we introduce the following concept.
@@
We say that $\mathscr{E}\leqslant\Omega^{\Omega}$ satisfies the finite retraction property if for every finite subset $A\subseteq\Omega$, there exists an element $h\in\mathscr{E}$ such that
@@
For instance, $\mathscr{E}=\Omega^{\Omega}$ clearly satisfies the finite retraction property.
@@
If $\mathscr{E}\leqslant\Omega^{\Omega}$ satisfies the finite retraction property, then every continuous representation of $\mathscr{E}$ is uniformly continuous.
@@
The order-preserving transformation monoid $\mathscr{E}=\operatorname{End}(\mathbb{Q},\leqslant)$ does not satisfy the finite retraction property. Indeed, for any finite subset $A\subseteq\mathbb{Q}$ with at least two points, there is no endomorphism $h\in\mathscr{E}=\operatorname{End}(\mathbb{Q},\leqslant)$ such that $h|_{A}=\operatorname{id}_{A}$ and $\mathrm{im}(h)\subseteq A$, because any order-preserving map from a dense linear order into a finite set must be locally constant, and thus cannot fix a set with more than one point. Consequently, $\mathscr{E}$ admits continuous representations that are not uniformly continuous.
@@ 7.6. Application III
For the third application, we let $\mathscr{E}\leqslant\Omega^{\Omega}$ and assume that the associated category $\mathscr{C}$} of finite subsets satisfies the following condition
@@
    every inclusion $\iota_{A,B}:A\hookrightarrow B$ admits a retraction given by a surjective map $p:B\twoheadrightarrow A$.
It is easy to see that the five categories ${\mathrm{OA}},{\mathrm{BA}},{\mathrm{CA}},{\mathrm{SA}},{\mathrm{FA}}$ all satisfy this condition.
@@
Let $\mathscr{E}\leqslant\Omega^{\Omega}$, and suppose that the associated category $\mathscr{C}$} of finite subsets satisfies condition $(\sharp)$. Then the realization functor $\mathcal{R}$ gives an equivalence
@@
Let $A$ be a finite subset and let $S\in J(A)$ be a covering sieve on ${\mathscr{C}}^{{\mathrm{op}}}$, which is the same as a left ideal on $\mathscr{C}$}. By the description of the Grothendieck topology $J$, the sieve $S$ contains an inclusion $\iota_{A,B}:A\hookrightarrow B$. By condition $(\sharp)$, there is a surjective map $p:B\twoheadrightarrow A$ such that
@@
Let $(\Omega,\mathcal{R}_{G})$ be a highly homogeneous relational structure. Without loss of generality, we may assume $\Omega=\mathbb{Q}$, on which the five highly homogeneous relations have been explicitly described in [2, Section 13]. Let $\mathscr{E}$ be the monoid of all self-maps of $\mathbb{Q}$ that preserve the highly homogeneous relation on $\Omega$. Then the equivalence described in Corollary 7.20 provides a convenient framework to study uniformly continuous representations of $\mathscr{E}$.
@@
Let $k$ be a commutative ring and fix $n\in\mathbb{Z}_{+}$. Let $G_{n}$ be the automorphism group of the finite subset $[n]$. Then under the equivalence in Corollary 7.20, we have the following correspondences:
@@
If furthermore $k$ is a field and $\lambda\in\operatorname{Irr}(G_{n})$, then
$(k\mathscr{M}_{n}/k\mathscr{N}_{n})\otimes_{kG_{n}}kG_{n}e_{\lambda}\,\longleftrightarrow\,\Delta_{n,\lambda},\quad(k\mathscr{M}_{n}/k\mathscr{N}_{n})\otimes_{kG_{n}}\lambda\,\longleftrightarrow\,\overline{\Delta}_{n,\lambda}.$
@@
One can then use the above result to obtain an explicit construction of irreducible uniformly continuous $k\mathscr{E}$-modules when $k$ is a field.
@@
Let $k$ be a field. The isomorphism classes of irreducible uniformly continuous $k\mathscr{E}$-modules are parameterized by the set of pairs
@@
Specifically, let $L_{n,\lambda}=(k\mathscr{M}_{n}/k\mathscr{N}_{n})\otimes_{kG_{n}}\lambda$. Then irreducible $k\mathscr{E}$-modules are described as follows:
@@
    If $\lambda$ is regular, the corresponding irreducible $k\mathscr{E}$-module is isomorphic to $L_{n,\lambda}$.
@@
    If $\lambda$ is singular, the corresponding irreducible $k\mathscr{E}$-module is isomorphic to the top of $L_{n,\lambda}$.
@@ 7.7. Application IV
The proofs of Proposition 7.6, Theorem 7.11, and Corollary 7.20 apply in this setting with suitable modifications. The orbit relations on $\Omega$ induced by the action of $G=\mathrm{GL}(\Omega)$ encode linear dependence of tuples. Consequently, the relevant finite subsets are precisely the finite dimensional $\mathbb{F}_{q}$-subspaces of $\Omega$. It follows that the category constructed from restrictions of elements of $\mathscr{E}$ is precisely ${\mathrm{VA}}_{q}$. Moreover, if $f,g\in\mathscr{E}$ coincide on a finite subset $A\subseteq\Omega$, then they coincide on the $\mathbb{F}_{q}$-linear span $\langle A\rangle$, since $f$ and $g$ are linear maps. Thus uniformly continuous $k\mathscr{E}$-modules can be defined via finite dimensional subspaces. Putting these observations together, we obtain the first equivalence in the first line.